OpenStudy (anonymous):

Find the axis of symmetry of the graph of the function. f(x) = 2x2 - 4x

OpenStudy (anonymous):

\[f(x)=2x^2-4x\\ y=2(x-1)^2 -(2)(1)\\ y=2(x-1)^2-2\\\\ y=a(x-h)^2+k\\ Focus=(1,- \frac{5}{8} )\\ y=-2-\frac{5}{8} \\ Directrix: y=- \frac{17}{8} \\ Axis of Symmetry: x=1 \]

OpenStudy (anonymous):

Thank You:)

OpenStudy (anonymous):

formula for axis of symmetry is: \[ -\frac{ b }{ 2 }\]